# NOLO: Navigate Only Look Once

### v1

@@ 4.1 Experimental Setups
Configurations. We apply our algorithm to open-source simulation-to-real platform RoboTHOR [83] and Habitat [84], which serve as realistic embodied navigation testbeds. For each navigation task in RoboTHOR and Habitat, we initialize a mobile robot with a random position, orientation, and a given goal image extracted from the context video indicating a specific goal object. At each timestep, the agent, mounted with only one RGB camera of $224\times 224$ resolution, receives an egocentric image and can execute an action from a fixed action space $[{\mathrm{MoveForward,TurnLeft,TurnRight,STOP}}]$. The minimum units of rotation and forward movement are 30°and 0.25 meters respectively. When the goal object is observed and the distance between the goal object and the agent is within 1.0 meter, the agent succeeds, otherwise it fails when the maximum time step is reached, which is 500 in all our experiments.
Task Division. In RoboTHOR, there are 15 different rooms and 5 possible layouts for each room, in total 75 different scenes. Each room has a different topology from other rooms, and the location of objects in each layout is different from other layouts. In order to test the in-context learning ability in different levels, we divided the training and test scenes as follows. For the training set, we select 12 rooms, and 4 layouts for each room. For the unseen layout testing set, we select the remaining 1 layout of the 12 training set rooms, in total 12 scenes. For the unseen room testing set, we select the remaining 3 rooms and all their layouts, in total 15 scenes. In Habitat, we divide all 90 building-scale scenes from Matterport3D (MP3D)  [85] into $70$ training scenes and $20$ testing scenes to test the generalization abilities. See Appendix C.1 for details.
Dataset Construction. We use a rule-based policy to collect the video dataset. The policy keeps moving forward until colliding with an obstacle and then randomly turns to get rid of it. Simultaneously, the agent keeps recording an egocentric traversal video till reaching the maximum 900 steps, i.e., a 30-second video clip in each scene. After that, we employ a pretrained object detector Detic [86] to detect all the objects occurred in the video. The collected egocentric traversal videos and detected goal frames are used for policy training and evaluation.
@@
For evaluation, we consider all detected goals in a video, and sample 10 robot start positions and orientations for each goal. Each start-goal pair becomes an evaluation task. We conduct 3 runs with different random seeds and quantitatively assess the performance using the following metrics:
  • Success Rate (SR): The percentage of tasks in which the agent reached the goal object.
  • Success weighted by normalized inverse Path Length (SPL[34]: Measured by the average number of steps taken to reach the goal object compared to the shortest possible path. Let $S_{i}$ be a binary indicator of success, $p_{i}$ be the length of the path actually taken by the agent and $\ell_{i}$ be the shortest path distance from the agent’s starting position to the goal in episode $i$, SPL is given as follows:
    $\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{\ell_{i}}{\max\left(p_{i},\ell_{i}\right)}.$ (8)
  • Trajectory Length (TL): The average number of steps for all trajectories.
  • Navigation Error (NE): The average distance between the final position of the agent to the goal object.
For visual clarity, we mainly plot the two most important metrics SR and SPL for performance comparison and leave TL and NE in Tables 1 and Appendix D. And we only highlight the metrics of NOLO in all our plots.
Baselines. It is worth noting that NOLO is the first method to incorporate context video into policy to do in-context navigation tasks. Therefore, we compare NOLO with random policy to indicate that NOLO can learn useful navigation policy from context videos. To prove the in-context generalization ability of NOLO, we additionally compare NOLO with two large multimodal models (LMM) GPT-4o and Video-LLaVA [87], which also exhibit remarkable zero-shot multimodal understanding and generation capabilities. To adapt these LMMs to the video navigation setting, a queue of size $K$ is allocated for maintaining context video frames, current observation, and goal image. Contents in the queue are concatenated as the input of LMMs. We design a consistent prompt to guide LMMs to adapt the navigation task, understand the context, and consequently output language responses both in RoboTHOR and Habitat, and then extract actions from the received language responses. See Appendix C.2 for more details.
@@ 4.2 RoboTHOR
((a)) Unseen Layout-SR
((b)) Unseen Layout-SPL
((c)) Unseen Room-SR
((d)) Unseen Room-SPL
Figure 4: Average SR and SPL in RoboTHOR seen rooms and unseen rooms.
We average the metrics of video navigation tasks in all testing scenes and showcase the mean results in Figure 4. Figure 4(a) and 4(b) show the performance in the unseen layout testing set. Figure 4(c) and 4(d) show the evaluation results in the unseen room testing set. NOLO demonstrates overwhelming advantages over baselines, including random and LMM baselines in both testing sets, which proves its strong generalization abilities to new scenes in different generalization levels. Considering an unseen room also means an unseen layout, it is a harder testing set than the unseen layout testing set in terms of the generalization ability. Therefore, NOLO displays slightly better navigation performance in the unseen layout testing set (71.92% SR and 29.26% SPL) than the unseen room testing set (70.48% SR and 27.74% SPL). GPT-4o outperforms random policy in most cases, while Video-LLaVA underperforms random policy in most cases. This is because the model size of Video-LLaVA is smaller than GPT-4o. The model size of NOLO is also smaller than GPT-4o, yet it shows better performance in the navigation problem, which again demonstrates the effectiveness of NOLO in training in-context navigation policy. See Appendix D for detailed results.
4.3 Habitat
Figure 5: Average SR in 20 Habitat testing scenes.
Figure 6: Average SPL in 20 Habitat testing scenes.
We evaluate NOLO similarly in Habitat to demonstrate the effectiveness of NOLO in larger-scale scenes. As illustrated in Figure 6 and 6, NOLO still outperforms other baselines in terms of SR and achieves slightly better performance in terms of SPL. Compared with the metrics of NOLO in RoboTHOR, the average SPL and SR are lower, and the average NE and TL (See Appendix D) are higher. This is mostly because scenes in Habitat are significantly larger than those in RoboTHOR, resulting in more challenging navigation tasks across multiple rooms and even floors and requiring longer traversal distances. However, we observe that NOLO, trained from 30-second video clips, is still capable of acquiring many necessary sub-skills like entering or exiting a room, avoiding obstacles, and exploring. See Appendix D for detailed results.
4.4 Cross-Domain Evaluation
To pursue the ultimate goal of video navigation, we evaluate NOLO across RoboTHOR and Habitat, even if the visual inputs may suffer from a severe domain gap. Specifically, NOLO trained from 48 RoboTHOR scenes is evaluated in 20 Habitat scenes (R2H), and NOLO trained from 70 Habitat scenes is evaluated in 27 RoboTHOR scenes (H2R). For comparison, we also plot the average in-domain results (R2R and H2H). We average all the results on testing scenes in Figure 7, from which we surprisingly discover that R2H and H2R show competitive performance compared to R2R and H2H. This is because the pretrained image encoder in NOLO provides a general representation of indoor scenes, and the temporal coherence loss (Equation (7)) also leads to a general representation reflecting the natural temporal ordering, which together lead to the generalization ability of NOLO across domains.
((a)) H2R-SR
((b)) H2R-SPL
((c)) R2H-SR
((d)) R2H-SPL
Figure 7: Average SR and SPL of cross domain evaluation results.
4.5 Ablations
@@
NOLO 71.12 $\pm$ 0.85 28.41 $\pm$ 0.72 244.21 $\pm$ 3.80 1.84 $\pm$ 0.04 43.65 $\pm$ 1.09 20.77 $\pm$ 0.46 347.57 $\pm$ 3.32 5.17 $\pm$ 0.09
NOLO-C 67.78 $\pm$ 0.60 27.70 $\pm$ 1.35 259.65 $\pm$ 0.53 1.98 $\pm$ 0.02 33.58 $\pm$ 0.92 17.41 $\pm$ 0.43 381.26 $\pm$ 3.23 6.58 $\pm$ 0.26
NOLO-T 69.03 $\pm$ 0.87 28.15 $\pm$ 0.09 252.84 $\pm$ 6.33 1.93 $\pm$ 0.04 37.48 $\pm$ 1.01 18.51 $\pm$ 0.71 369.85 $\pm$ 2.72 6.70 $\pm$ 0.07
NOLO(M) 66.15 $\pm$ 0.37 26.66 $\pm$ 1.61 267.95 $\pm$ 1.96 2.02 $\pm$ 0.00 36.73 $\pm$ 1.89 18.17 $\pm$ 0.31 373.92 $\pm$ 3.46 6.69 $\pm$ 0.25
@@
Does the context video provide cues for navigation? Yes. We conduct an ablation study by masking the context. We denote the ablation variant as NOLO-C, which is evaluated in RoboTHOR and Habitat with experimental configurations unchanged. We can see from Table 1 that a mean performance drop of NOLO-C is perceived both in RoboTHOR and Habitat compared with NOLO due to lack of context information.
@@
Does temporal coherence loss contribute to better visual representation and performance? Yes. We examine the effect of the temporal coherence loss (Equation (7)) in improving performance and representation. An ablation variant, named NOLO-T, is conducted by removing the temporal coherence loss. The results are shown in Table 1. The performance drops slightly in RoboTHOR while more obviously in Habitat. To figure out the underlying reasons for their discrepancy in performance, we further compare the learned visual representation of NOLO and NOLO-T. We randomly select a context video in RoboTHOR and utilize t-SNE projection to visualize their embedding sequences encoded by $E_{\zeta}(f_{t})$ in Figure 8(b). The embeddings learned by NOLO-T exhibit more irregular stochasticity. In contrast, NOLO learns a better-structured embedding, which demonstrates the effectiveness of the temporal coherence loss.Detailed results refer to Appendix D.
((a)) Visualization of saliency maps about context frames with respect to goal frames in RoboTHOR and Habitat.
((b)) Visualization of context embedding sequences of a sampled video from RoboTHOR.
Figure 8: We compare the saliency maps between NOLO and NOLO-C and context embedding sequences between NOLO and NOLO-T.
@@
Can NOLO incorporate other off-the-shelf modules to decode actions from videos? Yes. Key points matching methods like SuperGlue [88] can also be modified to extract actions from adjacent video frames. We follow a similar pipeline to decode actions via matching key points between adjacent video frames and then conduct BCQ. The new baseline, entitled NOLO(M), is evaluated under the same settings as NOLO. By referring to true action labels, we measure the average accuracy of action decoding of NOLO(M) and NOLO (80.43% vs 92.44%), which means the actions decoded from key point matching are less accurate than GMFlow. From Table 1, NOLO(M) exhibits around 5% decay in average success rate in RoboTHOR and 7% in Habitat. Therefore, to achieve more satisfactory in-context generalization, we choose GMFlow as the action decoder for NOLO. However, as a two-stage general framework for in-context learning from videos, NOLO is not limited to any kind of pseudo action labeling methods and is capable of incorporating more advanced modules for better performance.
Does NOLO understand the relation between goal and context video? Yes. To demonstrate the relation between goal and context video learned in NOLO, we visualize the Grad-CAM [89] saliency map of the visual encoder to showcase the effect of goal on video context. As demonstrated in Figure 8(a), when the goal image is given, NOLO precisely captures the important semantic cues in the goal image and wisely pays more attention to the related regions in the context frame, which means it builds connections between goal and context video. On the contrary, when the goal image is not given, NOLO cannot concentrate on the object specified in the goal image due to a lack of context guidance. Such a result indicates that NOLO can understand the goal image and capture relevant information in the context video.
@@ 5 Conclusion and Limitaiton
In this paper, we propose and investigate the video navigation problem, where the goal is to train an in-context navigation policy to find objects occurred in the context video, and the training of this policy has to be done in an offline manner purely from videos. To solve this problem, we propose our algorithm, NOLO, to train a generalizable in-context navigation policy taking a context video as input, such that it can generalize to a new scene without finetuning or re-training by understanding the context video of the new scene. In NOLO, we first extract pseudo actions from adjacent video frames using optical flow, and then use an offline reinforcement learning method, BCQ, to train the in-context navigation policy. We also propose a temporal coherence loss to learn representation consistent with the natural temporal ordering. We conduct experiments on Robothor and Habitat and demonstrate the generalization ability of NOLO in different levels. We also verify the effect of several key components of NOLO in ablation studies, which leads to a better understanding of how NOLO works.
It is important to acknowledge the limitations of NOLO, which currently can only accurately extract movement actions from two adjacent frames. Methods for decoding more complicated actions from multiple video frames are worth exploring. Additionally, in future work, we plan to pretrain NOLO on larger-scale scene datasets to verify its in-context generalization ability across more diverse scenes.
Appendix A Model Architecture
@@
Appendix B Methodology Details
B.1 Pseudo Action Labeling
We deploy optical flow and filter dominant vectors with the help of the flow map. Theoretically, discretizing filtered vectors into several action categories can be achieved by any unsupervised pattern recognition methods, such as clustering. In practice, we use a simple threshold-based heuristic. We observe that the feature of moving forward is less prominent than turning left or right. Consequently, we focus more on the horizontal displacement of the pixels in the flow map. If the horizontal component of horizontal displacement exceeds a threshold $\tau_{x}$ and the vertical component stays within a threshold $\tau_{y}$ then we classify it as a turning action, or otherwise it is considered as a forwarding action.
B.2 Policy Inference

### v2

@@ 4.1 Experimental Setups
Configurations. We apply our algorithm to realistic embodied navigation testbeds RoboTHOR [20] and Habitat [71]. For each navigation task in RoboTHOR and Habitat, we initialize a mobile robot with a random position, orientation, and a given goal image extracted from the context video indicating a specific goal object. At each timestep, the agent, mounted with only one RGB camera of $224\times 224$ resolution, receives an egocentric image and can execute a discrete action from a fixed action space $[{\mathrm{MoveForward,TurnLeft,TurnRight,STOP}}]$. The minimum units of rotation and forward movement are 30° and 0.25 meters respectively. When the goal object is observed and the distance between the goal object and the agent is within 1m, the agent succeeds, otherwise it fails when the maximum step is reached, which is 500 in all experiments.
Task Division. In RoboTHOR, there are 15 different rooms and 5 possible layouts for each room, in total 75 different scenes. Each room has a different topology from other rooms, and the location of objects in each layout is different from other layouts. In order to test the in-context learning ability at different levels, we divide the training and test scenes as follows. For the training set, we select 12 rooms, and 4 layouts for each room. For the unseen layout testing set, we select the remaining 1 layout of the 12 training set rooms, in total 12 scenes. For the unseen room testing set, we select the remaining 3 rooms and all their layouts, in total 15 scenes. In Habitat, we divide all 90 building-scale scenes from Matterport3D (MP3D)  [12] into $70$ training scenes and $20$ testing scenes to test the generalization abilities. See Appendix 7.1 for details.
Dataset Construction. We use a fixed non-explorative rule-based policy to collect the video dataset. The policy keeps moving forward until colliding with an obstacle and then randomly turns to get rid of it. Simultaneously, the agent keeps recording an egocentric RGB sequence till reaching the maximum 900 steps, i.e. a 30-second video clip in each scene. After that, we employ a pretrained object detector Detic [97] to detect all the objects occurred in the video. To get a set for goal frames, we additionally roll out in the simulation to collect goal images of each target object from different views.
@@
For evaluation, the agent is arranged to navigate to each target object detected in a context video. Each navigation task repeats 10 times.
We conduct 3 runs with different random seeds and quantitatively assess the performance using metrics including success rate (SR), success weighted by normalized inverse path length (SPL), trajectory length (TL), and navigation error (NE) following a recent work [93].
Baselines. NOLO stands as the first method to incorporate context videos directly into policy for in-context navigation tasks. To address in-context generalization capabilities, we compare NOLO with a random exploration policy, as well as two large multimodal models (LMMs), GPT-4 and Video-LLaVA which also exhibit remarkable zero-shot multimodal understanding and generation abilities, in a training-free manner. To adapt the LMMs for video-based navigation, we use a queue of size $K$ to store context video frames, current observations, and goal images, which are concatenated as LMM input. A unified prompt guides the LMMs to understand the context and navigate effectively in both RoboTHOR and Habitat, extracting actions from their language responses. See Appendix 7.2 for details.
In addition, we further compare NOLO with traditional visual navigation methods without context video input, specifically, an online method VGM [46], which builds a memory graph based on unsupervised image representations from navigation history to guide actions, and an offline method ZSON [60], which encodes goal images into a multimodal, semantic embedding space to enable scalable training of SemanticNav agents in unannotated 3D environments. Besides, we compare NOLO with a recent work AVDC [43], which infers actions from video predictions while incorporating extra depth information. Our experimental results show that AVDC underperforms NOLO even with additional depth images.
@@ 4.2 RoboTHOR
We average the metrics of video navigation tasks in all testing scenes and showcase the mean results in Tab. 1. NOLO demonstrates overwhelming advantages over baselines, including random, LMM baselines, and visual navigation baselines in both testing sets, which proves its strong generalization abilities to new scenes in different generalization levels. The set of unseen rooms is naturally a harder testing set than the unseen layouts in terms of generalization ability. Therefore, NOLO displays slightly better navigation performance in the unseen layout testing set (71.92% SR, 29.26% SPL) compared to the unseen room testing set (70.48% SR, 27.74% SPL). GPT-4o outperforms the random exploration policy across most scenarios, while Video-LLaVA generally performs below the random exploration policy due to its smaller model size relative to GPT-4o. NOLO is also much more parameter-efficient than GPT-4o. However, it achieves better navigation results, underscoring its effectiveness in training in-context navigation policies. Overall, visual navigation baselines tend to surpass both random and LMM baselines because of additional training. However, AVDC exhibits instability in unseen environments due to its reliance on a diffusion model trained on a limited set of egocentric video frames, which leads to inconsistency in video prediction. While VGM’s memory graph and ZSON’s multimodal semantic embeddings can provide some guidance, they still fall short of NOLO in terms of success rate and path efficiency. We analyzed that VGM’s implicit graph struggles to effectively represent topological structures with limited data coverage, and for ZSON, there is a non-negligible distribution discrepancy between CLIP [67] pre-training data and our egocentric videos. See Figs. 9 and 10 in Appendix 9 for detailed results.
Table 1: Performance comparison of success rates (SR), success path lengths (SPL), trajectory lengths (TL), and navigation errors (NE) across different methods and environments. The blue part shows some training-free methods and the green part shows visual navigation approaches which require training.
Method Robothor Habitat
Unseen Layout Unseen Room Unseen Room
[…]
Random 26.00 13.27 394.77 2.93 23.10 12.02 405.87 2.89 24.53 10.03 402.02 4.97
GPT-4o 31.97 17.18 363.23 3.03 31.69 16.49 364.63 2.89 35.16 20.39 345.84 4.70
Video-LLaVA 20.07 8.31 381.92 3.08 17.51 6.58 385.72 3.32 14.55 9.02 384.53 4.71
AVDC 32.53 9.43 430.77 2.15 31.14 11.57 424.56 2.13 27.29 12.24 418.12 4.30
VGM 47.04 23.90 376.92 2.44 40.51 19.99 400.90 2.38 34.34 17.30 393.92 4.19
ZSON 37.83 21.31 349.31 2.23 35.78 18.81 372.68 2.37 32.83 19.86 395.94 4.32
NOLO (ours) 71.92 29.26 238.92 1.79 70.48 27.74 248.44 1.87 43.65 20.77 347.57 3.67
(a) H2R-SR
(b) H2R-SPL
(c) R2H-SR
(d) R2H-SPL
Figure 4: Average SR and SPL of cross domain evaluation. H2R means training in Habitat and testing in RoboTHOR, and vice versa.
@@
NOLO 71.20 $\pm$ 0.85 28.50 $\pm$ 0.72 243.68 $\pm$ 3.80 1.83 $\pm$ 0.04 43.65 $\pm$ 1.09 20.77 $\pm$ 0.46 347.57 $\pm$ 3.32 3.67 $\pm$ 0.09
NOLO-C 67.78 $\pm$ 0.60 27.70 $\pm$ 1.35 259.65 $\pm$ 0.53 1.98 $\pm$ 0.02 33.58 $\pm$ 0.92 17.41 $\pm$ 0.43 381.26 $\pm$ 3.23 4.58 $\pm$ 0.26
NOLO-T 69.03 $\pm$ 0.87 28.15 $\pm$ 0.09 252.84 $\pm$ 6.33 1.93 $\pm$ 0.04 37.48 $\pm$ 1.01 18.51 $\pm$ 0.71 369.85 $\pm$ 2.72 4.72 $\pm$ 0.07
NOLO(M) 66.15 $\pm$ 0.37 26.66 $\pm$ 1.61 267.95 $\pm$ 1.96 2.02 $\pm$ 0.00 36.73 $\pm$ 1.89 18.17 $\pm$ 0.31 373.92 $\pm$ 3.46 4.69 $\pm$ 0.25
Figure 5: Visualizations of six video navigation tasks during in-context policy deployment. The leftmost bird-eye-view topological maps depicts the robot positions corresponding to the key observations to the right, and the second column displays the goal images.
4.3 Habitat
We evaluate NOLO similarly in Habitat to demonstrate the effectiveness of NOLO in larger-scale scenes. As illustrated in Tab. 1 , NOLO still outperforms baselines in terms of SR and achieves slightly better performance in terms of SPL. Compared with the metrics of NOLO in RoboTHOR, the average SPL and SR are lower, and the average NE and TL are higher. This is mostly because scenes in Habitat are significantly larger than those in RoboTHOR, resulting in more challenging navigation tasks across multiple rooms and even floors and requiring longer traversal distances. However, we observe that NOLO is still capable of acquiring many necessary sub-skills like entering or exiting a room, avoiding obstacles, and exploring. See Appendix 9 for detailed results.
4.4 Cross-Domain Evaluation
To pursue the ultimate goal of video navigation, we evaluate NOLO across RoboTHOR and Habitat, even if the visual inputs may suffer from a severe domain gap. Specifically, NOLO trained from 48 RoboTHOR scenes is evaluated in 20 Habitat scenes (R2H), and NOLO trained from 70 Habitat scenes is evaluated in 27 RoboTHOR scenes (H2R). For comparison, we also plot the average in-domain results (R2R and H2H). We average all the results on testing scenes in Fig. 4, from which we surprisingly discover that R2H and H2R show competitive performance compared to R2R and H2H. This is because the pretrained image encoder in NOLO provides a general representation of indoor scenes, and the temporal coherence loss in Eq. 5 also leads to a general representation reflecting the natural temporal ordering, which together lead to the generalization ability of NOLO across domains.
4.5 Ablation Studies
@@
Does the context video provide informative cues for navigation? Yes. We conduct an ablation study denoted as NOLO-C, in which we evaluate NOLO without context videos in RoboTHOR and Habitat with other configurations unchanged. From Tab. 2, a mean performance drop of NOLO-C is perceived both in RoboTHOR and Habitat compared with NOLO due to lack of context information.
@@
Does temporal coherence loss contribute to better visual representation and performance? Yes. We examine the effect of the temporal coherence loss in Eq. 5. An ablation variant, named NOLO-T, is conducted by removing the temporal coherence loss. The results are shown in Tab. 2. The performance drops slightly in RoboTHOR while more obviously in Habitat. To figure out the underlying reasons for their discrepancy in performance, we further compare the learned visual representation of NOLO and NOLO-T. We randomly select a context video in RoboTHOR and utilize t-SNE projection to visualize their embedding sequences encoded by $E_{\zeta}(f_{t})$ in Fig. 8 in Appendix 8. The embeddings learned by NOLO-T exhibit more irregular stochasticity. In contrast, NOLO learns a better-structured embedding, which demonstrates the effectiveness of the temporal coherence loss. See Appendix 9 for details.
@@
Can NOLO incorporate other off-the-shelf modules to decode actions from videos? Yes. We conduct an ablation, entitled NOLO(M), to decode actions from SuperGlue [70] via point matching instead of GMFlow [89]. We observe an accuracy drop in action decoding (from 92.44% to 80.43%). We then follow the same pipeline and report the performance in Tab. 2. NOLO(M) exhibits around 5% decay in average success rate in RoboTHOR and 7% in Habitat. Therefore, to achieve more satisfactory in-context generalization, we choose GMFlow as the action decoder for NOLO. However, as a general two-stage framework for in-context learning from videos, NOLO is not limited to any kind of pseudo action labeling methods and is capable of incorporating more advanced modules for better performance.
Does NOLO understand the relationship between goal and context video? Yes. We use Grad-CAM [73] saliency maps of the visual encoder to illustrate NOLO’s ability to connect the goal with the context video. As shown in Fig. 6, when a goal image is provided, NOLO accurately identifies key semantic cues in the goal and focuses on related areas in the context frame, indicating its capacity to link goal and context. Without the goal image, however, NOLO lacks this focus, underscoring its reliance on goal-driven context understanding. This result demonstrates NOLO’s ability to interpret the goal and capture relevant information in the context video.
4.6 Real-World Experiments
We constructed a real-world maze environment, as shown in Fig. 7. We employed the Unitree Go2 robot, equipped with a RealSense D435i camera to capture RGB observations. In the maze environment, the robot’s actions are discretized into moving forward by 15 cm or turning 30 degrees at each step. Six target objects–a book, box, cola can, cup, glue, and novel–are placed in the maze. Images of each target are taken from random nearby perspectives. A fixed, non-exploratory, rule-based policy was employed to collect two 600-step video clips in the maze, one for training and the other as context video during policy deployment. For evaluation, each episode consists of 100 steps, with a success radius set at 50 cm. Fig. 5 illustrates key frames from successful trajectories during in-context policy deployment. The robot demonstrated impressive capabilities in navigating straight passages, executing turns at corners, and ultimately reaching the goals by comprehending the dynamics and topological structure inferred from the unseen context video and its relationship to the goal images.
@@ 5 Conclusion and Limitaiton
Figure 6: Comparision of saliency maps between NOLO and NOLO-C about context frames with respect to goal frames in RoboTHOR and Habitat.
Figure 7: Setups of real-world maze. The robotic dog mounted with a head camera acts as the navigation agent and six annotated objects are goals.
@@
Supplementary Material
6 Methodology Details
6.1 Pseudo Action Labeling
We deploy optical flow and filter dominant vectors with the help of the flow map. Theoretically, discretizing filtered vectors into several action categories can be achieved by any unsupervised pattern recognition methods, such as clustering. In practice, we use a simple threshold-based heuristic. We observe that the feature of moving forward is less prominent than turning left or right. Consequently, we focus more on the horizontal displacement of the pixels in the flow map. If the horizontal component of horizontal displacement exceeds a threshold $\tau_{x}$ and the vertical component stays within a threshold $\tau_{y}$ then we classify it as a turning action, or it is considered as a forwarding action.
Occasionally, we notice that the displacement of pixels observed across frames manifests irregular patterns. To mitigate this issue, for each pair of adjacent frames $(f_{t},f_{t+1})$, we filter dominant vectors $\nu_{t}$ that exhibit gradient amplitudes within the upper decile, which are subsequently discretized into actions derived from the action space, making it possible to trace the trajectory of the agent over time.
6.2 Policy Inference